In a group of 100 batteries, 30 are dead.
If you choose two batteries, what is the probability that they both will be dead?

Answer by checkley77(12569)

30/100*29/99=870/9,900=.088 OR 8.8% CHANCE OF SELECTING 2 DEAD BATTERIES.